Suppose that the functions s and t are defined for all real numbers x as follows. 

s(x)=x-6 

t(x)= 3x2 


Write the expressions for (s – t)(x) and (s·t)(x) and evaluate (s+t)(-3). 

(s – t)(x) = _______ 

 (s·t)(x) = _______  

(s + t)(-3) = _______ 
Answer #1

(s-t)(x) = (x-6) - 3x2

(s*t)(x) = (x-6)*3x2 = 3x3 - 18x2

(s+t)(x) = (x-6) - 3x2

(s+t)(-3) = (-3-6) - 3(-3)2 = -9 -27 = -36